# Constraints on Long-Ranged Interactions Between Dark Matter and the Standard Model

### v1

@@ III.2 Effects of Dynamical Friction
For Coulomb-like scattering with a velocity-dependent momentum-transfer cross-section $\sigma_{T}=\sigma_{0}v^{-4}$ and $m_{*}\gg m_{\chi}$, the rate of heat transfer from stars to dark matter is given by [41, 42]
$\frac{d^{2}U_{\chi}}{dt\,dV}$ $=\frac{2\rho_{*}\rho_{\chi}\sigma_{0}(T_{*}-T_{\chi})}{\sqrt{2\pi}m_{*}^{2}v_{\rm th}^{3}}$ (16)
@@
$v_{\rm th}=\sqrt{\frac{T_{*}}{m_{*}}+\frac{T_{\chi}}{m_{\chi}}}.$ (17)
@@
$\frac{d^{2}U_{\chi}}{dt\,dV}$ $=4\sqrt{2\pi}G^{2}\rho_{*}\rho_{\chi}\frac{m_{*}\sigma_{*}^{2}-m_{\chi}\sigma_{\chi}^{2}}{(\sigma_{*}^{2}+\sigma_{\chi}^{2})^{3/2}}\ln\Lambda_{G}$ (18)
@@
$b_{\rm max}$ $=R_{*}$ (20)
$b_{\rm min}$ $=\frac{Gm_{*}}{\sigma_{*}^{2}+\sigma_{\chi}^{2}}$ (21)
@@
Computing the total heat transfer rate from stars to dark matter would require us to integrate Eq. (16) over the UFD’s volume. While this could be performed numerically, we can obtain approximate analytic results for this total heating rate by taking advantage of the small variation in the temperatures of the stars and dark matter at radii $r\lesssim R_{*}$, where the majority of heat transfer happens; while not exact, this is sufficient for the order-one estimates in this work, and is numerically less expensive than the numerical alternative. In this approximation, the total heat transfer becomes
$\begin{split}\frac{dU_{\chi}}{dt}\approx\,&4\sqrt{2\pi}G^{2}\frac{M_{*,{\rm tot}}M_{\chi,{\rm tot}}}{R_{*}^{3}}\mathcal{F}\\ &\times\frac{m_{*}\sigma_{*}(R_{*})^{2}-m_{\chi}\sigma_{\chi}(R_{*})^{2}}{(\sigma_{*}(R_{*})^{2}+\sigma_{\chi}(R_{*})^{2})^{3/2}}\ln\Lambda_{G}(R_{*})\end{split}$ (22)
@@
$\mathcal{F}$ $=\frac{2}{\pi^{5/2}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}0&\frac{1}{2}&1\\[2.0pt] \frac{3}{2}&2&\frac{5}{2}\end{matrix}\right.\right)$ (24)
@@
Figure 2: The stellar scale radius of Segue I at various times in the past, assuming dynamical friction from gravity alone and three different halo profiles: Dehnen spheres with $\gamma=0$ and $\gamma=1$ and an NFW profile. Also shown are the (constant) dark matter scale radii of each profile, obtained by combining the simulated total mass and the observed stellar velocity dispersion. The maximum times in the past for each line are chosen so that the stellar scale radius would have been greater than the dark matter scale radius, since the unphysical nature of such a galaxy is the basis for the constraints we set (see Sections III.3 and III.4); while the times required for this are unphysical in the gravity-only case, the introduction of new forces is largely equivalent to rescaling the time, leading to significant evolution over physically-meaningful timescales.
In the presence of new forces, Fig. 2 is modified for two reasons: First, there is a direct $\alpha_{\rm SD}^{2}$ dependence in the total heating rate (i.e. $G^{2}\to\alpha_{\rm SD}^{2}G^{2}$ in Eq. (22)), which accelerates dynamical friction by a factor of $\alpha_{\rm SD}^{2}$. (There is additionally a logarithmic dependence on $\alpha_{\rm SD}$ and on the new force range $\lambda$ through a modified logarithm (see Eqs. (47) and (52) in Appendix B), but this is only significant near the shortest ranges where our bound applies). If this quadratic contribution were the only modification to the heating rate, we could immediately set constraints on new forces based on a conservative assumption about galaxy formation: that the initial stellar scale radius be smaller than the dark matter scale radius, since this is where stars are expected to form. The maximum allowed $\alpha_{\rm SD}$ (for $\lambda>R_{*}$) would then be given by
@@
However, there is a second modification to the rate of dynamical friction caused by the introduction of a new dark matter-Standard Model force. A new force of this type also implies the existence of dark matter-dark matter and Standard Model-Standard Model forces, as discussed in Section II. While the small mass of Standard Model particles in UFDs makes the latter irrelevant, the introduction of a strong ($\alpha_{\rm DD}\gg 1$, for most of the parameter space we consider), long-range new force between dark matter particles necessarily significantly modifies the structure of the dark matter halo itself, as we discuss in Section III.3.
The exception to this is for dark matter with zero net charge (e.g. equal numbers of particles with opposite charges), in which case dynamical friction remains present (since it requires a net charge for stars, but not for dark matter) but most such modifications to the halo itself do not (since there are no large net forces when the dark matter is net-neutral). However, in this case, one must account for the effects of Debye screening and plasma instabilities. We consider these in Section III.4.
III.3 Dynamical Friction with Net-Charged Dark Matter
@@
Introducing a force between dark matter particles would modify the condition for hydrostatic equilibrium in the halo, Eq. (13), requiring a change in the total mass, scale radius (keeping the profile shape fixed), and/or temperature distribution of the halo in order to preserve equilibrium. Similarly, introducing a force between dark matter particles and Standard Model particles affects the condition for the stars’ hydrostatic equilibrium, Eq. (12). The latter, however, can be resolved only by adjusting the central dark matter density of the profile, since the stars’ temperature distribution is (roughly) fixed by their observed velocity dispersion. For any chosen profile shape (i.e. $\gamma=0$, $\gamma=1$, or NFW), there are two ways to change this central density: a change in the total mass of the halo, or a change in its scale radius (or a combination of the two).
@@
$\alpha_{\rm SD}^{\rm(eff)}$ $\sim\frac{\alpha_{\rm SD}}{1+(R_{\chi}/\lambda)^{2}}\qquad{\rm for~{}}\gamma=0$ (27a)
$\alpha_{\rm DD}^{\rm(eff)}$ $\sim\frac{\alpha_{\rm DD}}{1+(R_{\chi}/\lambda)^{2}}\qquad{\rm for~{}}\gamma=0$ (27b)
@@
$\alpha_{\rm SD}^{\rm(eff)}$ $\sim\frac{\alpha_{\rm SD}}{1+(R_{*}/\lambda)^{2}}\qquad{\rm for~{}NFW~{}or~{}}\gamma=1$ (28a)
$\alpha_{\rm DD}^{\rm(eff)}$ $\sim\frac{\alpha_{\rm DD}}{1+(R_{*}/\lambda)^{2}}\qquad{\rm for~{}NFW~{}or~{}}\gamma=1.$ (28b)
@@
We can now evaluate the changes to Segue I’s halo properties, and thus the changes to our dynamical friction calculation, in the presence of new forces. We begin with universally attractive forces. In this case, the density of dark matter in the halo must be reduced by a factor of $1+\alpha_{\rm SD}^{\rm(eff)}$, in order to maintain the observed stellar velocity dispersion. This in-turn reduces the dark matter velocity dispersion by a factor of $\sqrt{1+\alpha_{\rm SD}^{\rm(eff)}}$ (see Eq. (14b)), but there is a competing increase in $\chi$’s velocity caused by the new dark matter-dark matter force; this enhancement is given by $\sqrt{1+\alpha_{\rm DD}^{\rm(eff)}}$.
@@
At this point we see an additional complication: we wish to constrain $\alpha_{\rm SD}$, but our result depends on $\alpha_{\rm DD}$ as well. Since the heating rate Eqn. (22) is, for $T_{\chi}\ll T_{*}$, proportional to $M_{\chi}\sigma_{\chi}^{-3/2}$ until $\sigma_{\chi}=\sigma_{*}$ (after which it is independent of $\sigma_{\chi}$), the weakest constraint on $\alpha_{\rm SD}$ will arise when $\alpha_{\rm DD}$ is equal to the maximum value allowed by other observations, $\alpha_{\rm DD}^{\rm(max)}$. The time needed for the stellar scale radius to reach its present value, starting from the dark matter scale radius, is thus
@@
$\tau_{\rm eq}$ $\sim\tau_{\rm eq}^{G}\alpha_{\rm SD}^{-2}\frac{\ln\Lambda_{G}}{\ln\Lambda_{Y}}\times\begin{cases}\frac{\left(1+\alpha_{\rm DD}^{\rm(max)}\right)^{3/2}}{\sqrt{1+\alpha_{\rm SD}^{\rm(eff)}}}&\sigma_{\chi}\gtrsim\sigma_{*}\\ {\left(\frac{\sigma_{*,0}}{\sigma_{\chi,0}}\right)}^{3/2}\left(1+\alpha_{\rm SD}^{\rm(eff)}\right)&\sigma_{\chi}\lesssim\sigma_{*}\end{cases}$ (29)
@@
where $\Lambda_{G}$ is the Coulomb logarithm for gravity (Eqn. (19)) while $\Lambda_{Y}$ is the analogous logarithm for a Yukawa force (see Eq. (52) in Appendix B) and $\sigma_{*,0}$ ($\sigma_{\chi,0}$) are the velocity dispersions of the stars (of $\chi$) calculated assuming no new forces; $\Lambda_{Y}$ implicitly depends on $\alpha_{\rm SD}$ and $\lambda$. Note the presence of both $\alpha_{\rm SD}$ and $\alpha_{\rm SD}^{\rm(eff)}$ in Eq. (29): while the reduction of the dark matter density is suppressed for $R_{\chi}\gg\lambda$, dynamical friction is not.
@@
Requiring that $\tau_{\rm eq}\gtrsim 10$ Gyr thus gives us an implicit solution for our constraint. This is plotted, for each of the profile shapes we consider, in Figs. 3 and 4 (including two example dark matter masses in each, since the mass of $\chi$ affects the value of $\alpha_{\rm DD}^{\rm(max)}$). The two figures differ in their assumptions about the existing constraints on Standard Model-Standard Model forces; see Fig. 1 and the discussion in Section II. We discuss some features of these figures below, after handling the remaining cases.
@@
Figure 3: The upper bounds on the strength $\alpha_{\rm SD}$ of a new force as a function of its range $\lambda$, see Eq. (1) (with $q_{S}=q_{D}=1$). This figure covers the case of an attractive force between dark matter (DM) and either an attractive or repulsive force between dark matter and Standard Model (SM) particles, in the limit of zero equivalence principle (EP) violation arising from the SM-SM force. Shown are the bounds from combining constraints on DM-DM interactions with constraints on SM-SM interactions we present in Section II, as well as the bounds from dynamical friction in ultrafaint dwarf galaxies (UFDs) that we derive in Section III. The non-dynamical friction bounds are shown for two example DM masses $m_{\rm DM}$, along with a DM mass-independent bound (arising from the need for stars to stay bound to UFDs) that applies only if DM and SM particles repel one another. The dynamical friction bounds are likewise shown for both DM masses and for both the attractive and repulsive SM-DM cases, although the distinction between the latter two is small (except that the repulsive case is nonsensical beyond the UFD stellar binding line). Results are shown only for a $\gamma=0$ profile; no constraint can be obtained for the $\gamma=1$ or NFW profiles when the new force attracts dark matter to itself.
@@
Figure 4: Similar to Fig. 3 but here with EP violation in the SM-SM force. This figure shows the upper bounds on new forces that attract dark matter (DM) particles to one another and either attract or repel dark matter to/from Standard Model (SM) particles, in terms of their range $\lambda$ and their strength relative to gravity $\alpha_{\rm SD}$—see Eq. (1) (with $q_{S}=q_{D}=1$)—in the limit of maximal equivalence principle (EP) violation arising from the SM-SM force. Other than the change in bounds on the SM-SM force in the EP-violating case, and the addition of the torsion balance-based test of Milky Way center-directed equivalence principle violation in Ref. [25], this plot is identical to Fig. 3; see that figure’s caption for details.
@@
For forces that are attractive among $\chi$ particles but repulsive between $\chi$ and stars, $\alpha_{\rm SD}^{\rm(eff)}$ now increases the inferred halo mass rather than decreasing it; this changes a sign in Eq. (29) and eliminates the need to consider $\sigma_{\chi}\lesssim\sigma_{*}$:
$\tau_{\rm eq}$ $\sim\tau_{\rm eq}^{G}\alpha_{\rm SD}^{-2}\frac{\ln\Lambda_{G}}{\ln\Lambda_{Y}}\times\frac{\left(1+\alpha_{\rm DD}^{\rm(max)}\right)^{3/2}}{\sqrt{1-\alpha_{\rm SD}^{\rm(eff)}}}.$ (30)
@@
$\tau_{\rm eq}$ $\sim\tau_{\rm eq}^{G}\alpha_{\rm SD}^{-2}\frac{\ln\Lambda_{G}}{\ln\Lambda_{Y}}\times\begin{cases}\frac{\left(1-\frac{\alpha_{\rm SD}^{\rm(eff)}\alpha_{\rm SD}}{\alpha_{\rm SS}^{\rm(max)}}\right)^{3/2}}{\sqrt{1+\alpha_{\rm SD}^{\rm(eff)}}}&\sigma_{\chi}\gtrsim\sigma_{*}\\ {\left(\frac{\sigma_{*,0}}{\sigma_{\chi,0}}\right)}^{3/2}\left(1+\alpha_{\rm SD}^{\rm(eff)}\right)&\sigma_{\chi}\lesssim\sigma_{*}\end{cases}$ (31)
@@
$\tau_{\rm eq}$ $\sim\tau_{\rm eq}^{G}\alpha_{\rm SD}^{-2}\frac{\ln\Lambda_{G}}{\ln\Lambda_{Y}}\times\begin{cases}\frac{\left(1-\frac{\alpha_{\rm SD}^{\rm(eff)}\alpha_{\rm SD}}{\alpha_{\rm SS}^{\rm(max)}}\right)^{3/2}}{\sqrt{1-\alpha_{\rm SD}^{\rm(eff)}}}&\sigma_{\chi}\gtrsim\sigma_{*}\\ {\left(\frac{\sigma_{*,0}}{\sigma_{\chi,0}}\right)}^{3/2}\left(1-\alpha_{\rm SD}^{\rm(eff)}\right)&\sigma_{\chi}\lesssim\sigma_{*}\end{cases}$ (32)

### v2

@@ 3.2 Effects of Dynamical Friction
For Coulomb-like scattering with a velocity-dependent momentum-transfer cross-section $\sigma_{T}=\sigma_{0}v^{-4}$ and $m_{*}\gg m_{\chi}$, the rate of heat transfer from stars to dark matter is given by [45, 46, 47]
$\frac{d^{2}U_{\chi}}{dt\,dV}$ $=\frac{2\rho_{*}\rho_{\chi}\sigma_{0}(T_{*}-T_{\chi})}{\sqrt{2\pi}m_{*}^{2}v_{\rm th}^{3}}$ (3.10)
@@
$v_{\rm th}=\sqrt{\frac{T_{*}}{m_{*}}+\frac{T_{\chi}}{m_{\chi}}}.$ (3.11)
@@
$\frac{d^{2}U_{\chi}}{dt\,dV}$ $=4\sqrt{2\pi}G^{2}\rho_{*}\rho_{\chi}\frac{m_{*}\sigma_{*}^{2}-m_{\chi}\sigma_{\chi}^{2}}{(\sigma_{*}^{2}+\sigma_{\chi}^{2})^{3/2}}\ln\Lambda_{G}$ (3.12)
@@
$b_{\rm max}$ $=R_{*}$ (3.14)
$b_{\rm min}$ $=\frac{Gm_{*}}{\sigma_{*}^{2}+\sigma_{\chi}^{2}}$ (3.15)
@@
Computing the total heat transfer rate from stars to dark matter would require us to integrate Eq. (3.10) over the UFD’s volume. While this could be performed numerically, we can obtain approximate analytic results for this total heating rate by taking advantage of the small variation in the temperatures of the stars and dark matter at radii $r\lesssim R_{*}$, where the majority of heat transfer happens; while not exact, this is sufficient for the order-one estimates in this work, and is numerically less expensive than the numerical alternative. In this approximation, the total heat transfer becomes
$\begin{split}\frac{dU_{\chi}}{dt}\approx\,&4\sqrt{2\pi}G^{2}\frac{M_{*,{\rm tot}}M_{\chi,{\rm tot}}}{R_{*}^{3}}\mathcal{F}\frac{m_{*}\sigma_{*}(R_{*})^{2}-m_{\chi}\sigma_{\chi}(R_{*})^{2}}{(\sigma_{*}(R_{*})^{2}+\sigma_{\chi}(R_{*})^{2})^{3/2}}\ln\Lambda_{G}(R_{*})\end{split}$ (3.16)
@@
$\mathcal{F}$ $=\frac{2}{\pi^{5/2}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}0&\frac{1}{2}&1\\[2.0pt] \frac{3}{2}&2&\frac{5}{2}\end{matrix}\right.\right)$ (3.18)
@@
Figure 2: The stellar scale radius of Segue I at various times in the past, assuming dynamical friction from gravity alone and three different halo profiles: Dehnen spheres with $\gamma=0$ and $\gamma=1$ and an NFW profile. Also shown are the (constant) dark matter scale radii of each profile, obtained by combining the simulated total mass and the observed stellar velocity dispersion. The maximum times in the past for each line are chosen so that the stellar scale radius would have been greater than the dark matter scale radius, since the unphysical nature of such a galaxy is the basis for the constraints we set (see Sections 3.3 and 3.4); while the times required for this are unphysical in the gravity-only case, the introduction of new forces is largely equivalent to rescaling the time, leading to significant evolution over physically-meaningful timescales.
In the presence of new forces, Fig. 2 is modified for two reasons: First, there is a direct $\alpha_{\rm SD}^{2}$ dependence in the total heating rate (i.e. $G^{2}\to\alpha_{\rm SD}^{2}G^{2}$ in Eq. (3.16)), which accelerates dynamical friction by a factor of $\alpha_{\rm SD}^{2}$. (There is additionally a logarithmic dependence on $\alpha_{\rm SD}$ and on the new force range $\lambda$ through a modified logarithm (see Eqs. (B.6) and (B.11) in Appendix B), but this is only significant near the shortest ranges where our bound applies). If this quadratic contribution were the only modification to the heating rate, we could immediately set constraints on new forces based on a conservative assumption about galaxy formation: that the initial stellar scale radius be smaller than the dark matter scale radius, since this is where stars are expected to form. The maximum allowed $\alpha_{\rm SD}$ (for $\lambda>R_{*}$) would then be given by
@@
However, there is a second modification to the rate of dynamical friction caused by the introduction of a new dark matter-Standard Model force. A new force of this type also implies the existence of dark matter-dark matter and Standard Model-Standard Model forces, as discussed in Section 2. While the small mass of Standard Model particles in UFDs makes the latter irrelevant, the introduction of a strong ($\alpha_{\rm DD}\gg 1$, for most of the parameter space we consider), long-range new force between dark matter particles necessarily significantly modifies the structure of the dark matter halo itself, as we discuss in Section 3.3.
The exception to this is for dark matter with zero net charge (e.g. equal numbers of particles with opposite charges), in which case dynamical friction remains present (since it requires a net charge for stars, but not for dark matter) but most such modifications to the halo itself do not (since there are no large net forces when the dark matter is net-neutral). However, in this case, one must account for the effects of Debye screening and plasma instabilities. We consider these in Section 3.4.
3.3 Dynamical Friction with Net-Charged Dark Matter
@@
Introducing a force between dark matter particles would modify the condition for hydrostatic equilibrium in the halo, Eq. (3.7), requiring a change in the total mass, scale radius (keeping the profile shape fixed), and/or temperature distribution of the halo in order to preserve equilibrium. Similarly, introducing a force between dark matter particles and Standard Model particles affects the condition for the stars’ hydrostatic equilibrium, Eq. (3.6). The latter, however, can be resolved only by adjusting the central dark matter density of the profile, since the stars’ temperature distribution is (roughly) fixed by their observed velocity dispersion. For any chosen profile shape (i.e. $\gamma=0$, $\gamma=1$, or NFW), there are two ways to change this central density: a change in the total mass of the halo, or a change in its scale radius (or a combination of the two).
@@
$\alpha_{\rm SD}^{\rm(eff)}$ $\sim\frac{\alpha_{\rm SD}}{1+(R_{\chi}/\lambda)^{2}}\qquad{\rm for~{}}\gamma=0$ (3.21a)
$\alpha_{\rm DD}^{\rm(eff)}$ $\sim\frac{\alpha_{\rm DD}}{1+(R_{\chi}/\lambda)^{2}}\qquad{\rm for~{}}\gamma=0$ (3.21b)
@@
$\alpha_{\rm SD}^{\rm(eff)}$ $\sim\frac{\alpha_{\rm SD}}{1+(R_{*}/\lambda)^{2}}\qquad{\rm for~{}NFW~{}or~{}}\gamma=1$ (3.22a)
$\alpha_{\rm DD}^{\rm(eff)}$ $\sim\frac{\alpha_{\rm DD}}{1+(R_{*}/\lambda)^{2}}\qquad{\rm for~{}NFW~{}or~{}}\gamma=1.$ (3.22b)
@@
We can now evaluate the changes to Segue I’s halo properties, and thus the changes to our dynamical friction calculation, in the presence of new forces. We begin with universally attractive forces. In this case, the density of dark matter in the halo must be reduced by a factor of $1+\alpha_{\rm SD}^{\rm(eff)}$, in order to maintain the observed stellar velocity dispersion. This in-turn reduces the dark matter velocity dispersion by a factor of $\sqrt{1+\alpha_{\rm SD}^{\rm(eff)}}$ (see Eq. (3.8b)), but there is a competing increase in $\chi$’s velocity caused by the new dark matter-dark matter force; this enhancement is given by $\sqrt{1+\alpha_{\rm DD}^{\rm(eff)}}$.
@@
At this point we see an additional complication: we wish to constrain $\alpha_{\rm SD}$, but our result depends on $\alpha_{\rm DD}$ as well. Since the heating rate Eqn. (3.16) is, for $T_{\chi}\ll T_{*}$, proportional to $M_{\chi}\sigma_{\chi}^{-3/2}$ until $\sigma_{\chi}=\sigma_{*}$ (after which it is independent of $\sigma_{\chi}$), the weakest constraint on $\alpha_{\rm SD}$ will arise when $\alpha_{\rm DD}$ is equal to the maximum value allowed by other observations, $\alpha_{\rm DD}^{\rm(max)}$. The time needed for the stellar scale radius to reach its present value, starting from the dark matter scale radius, is thus
@@
$\tau_{\rm eq}$ $\sim\tau_{\rm eq}^{G}\alpha_{\rm SD}^{-2}\frac{\ln\Lambda_{G}}{\ln\Lambda_{Y}}\times\begin{cases}\frac{\left(1+\alpha_{\rm DD}^{\rm(max)}\right)^{3/2}}{\sqrt{1+\alpha_{\rm SD}^{\rm(eff)}}}&\sigma_{\chi}\gtrsim\sigma_{*}\\ {\left(\frac{\sigma_{*,0}}{\sigma_{\chi,0}}\right)}^{3/2}\left(1+\alpha_{\rm SD}^{\rm(eff)}\right)&\sigma_{\chi}\lesssim\sigma_{*}\end{cases}$ (3.23)
@@
where $\Lambda_{G}$ is the Coulomb logarithm for gravity (Eqn. (3.13)) while $\Lambda_{Y}$ is the analogous logarithm for a Yukawa force (see Eq. (B.11) in Appendix B) and $\sigma_{*,0}$ ($\sigma_{\chi,0}$) are the velocity dispersions of the stars (of $\chi$) calculated assuming no new forces; $\Lambda_{Y}$ implicitly depends on $\alpha_{\rm SD}$ and $\lambda$. Note the presence of both $\alpha_{\rm SD}$ and $\alpha_{\rm SD}^{\rm(eff)}$ in Eq. (3.23): while the reduction of the dark matter density is suppressed for $R_{\chi}\gg\lambda$, dynamical friction is not.
@@
Requiring that $\tau_{\rm eq}\gtrsim 10$ Gyr thus gives us an implicit solution for our constraint. This is plotted, for each of the profile shapes we consider, in Figs. 3 and 4 (including two example dark matter masses in each, since the mass of $\chi$ affects the value of $\alpha_{\rm DD}^{\rm(max)}$). The two figures differ in their assumptions about the existing constraints on Standard Model-Standard Model forces; see Fig. 1 and the discussion in Section 2. We discuss some features of these figures below, after handling the remaining cases.
@@
Figure 3: The upper bounds on the strength $\alpha_{\rm SD}$ of a new force as a function of its range $\lambda$, see Eq. (1.1) (with $q_{S}=q_{D}=1$). This figure covers the case of an attractive force between dark matter (DM) and either an attractive or repulsive force between dark matter and Standard Model (SM) particles, in the limit of zero equivalence principle (EP) violation arising from the SM-SM force. Shown are the bounds from combining constraints on DM-DM interactions with constraints on SM-SM interactions we present in Section 2, as well as the bounds from dynamical friction in ultrafaint dwarf galaxies (UFDs) that we derive in Section 3. The non-dynamical friction bounds are shown for two example DM masses $m_{\rm DM}$, along with a DM mass-independent bound (arising from the need for stars to stay bound to UFDs) that applies only if DM and SM particles repel one another. The dynamical friction bounds are likewise shown for both DM masses and for both the attractive and repulsive SM-DM cases, although the distinction between the latter two is small (except that the repulsive case is nonsensical beyond the UFD stellar binding line). Results are shown only for a $\gamma=0$ profile; no constraint can be obtained for the $\gamma=1$ or NFW profiles when the new force attracts dark matter to itself.
@@
Figure 4: Similar to Fig. 3 but here with EP violation in the SM-SM force. This figure shows the upper bounds on new forces that attract dark matter (DM) particles to one another and either attract or repel dark matter to/from Standard Model (SM) particles, in terms of their range $\lambda$ and their strength relative to gravity $\alpha_{\rm SD}$—see Eq. (1.1) (with $q_{S}=q_{D}=1$)—in the limit of maximal equivalence principle (EP) violation arising from the SM-SM force. Other than the change in bounds on the SM-SM force in the EP-violating case, and the addition of the torsion balance-based test of Milky Way center-directed equivalence principle violation in Ref. [27], this plot is identical to Fig. 3; see that figure’s caption for details.
@@
For forces that are attractive among $\chi$ particles but repulsive between $\chi$ and stars, $\alpha_{\rm SD}^{\rm(eff)}$ now increases the inferred halo mass rather than decreasing it; this changes a sign in Eq. (3.23) and eliminates the need to consider $\sigma_{\chi}\lesssim\sigma_{*}$:
$\tau_{\rm eq}$ $\sim\tau_{\rm eq}^{G}\alpha_{\rm SD}^{-2}\frac{\ln\Lambda_{G}}{\ln\Lambda_{Y}}\times\frac{\left(1+\alpha_{\rm DD}^{\rm(max)}\right)^{3/2}}{\sqrt{1-\alpha_{\rm SD}^{\rm(eff)}}}.$ (3.24)
@@
$\tau_{\rm eq}$ $\sim\tau_{\rm eq}^{G}\alpha_{\rm SD}^{-2}\frac{\ln\Lambda_{G}}{\ln\Lambda_{Y}}\times\begin{cases}\frac{\left(1-\frac{\alpha_{\rm SD}^{\rm(eff)}\alpha_{\rm SD}}{\alpha_{\rm SS}^{\rm(max)}}\right)^{3/2}}{\sqrt{1+\alpha_{\rm SD}^{\rm(eff)}}}&\sigma_{\chi}\gtrsim\sigma_{*}\\ {\left(\frac{\sigma_{*,0}}{\sigma_{\chi,0}}\right)}^{3/2}\left(1+\alpha_{\rm SD}^{\rm(eff)}\right)&\sigma_{\chi}\lesssim\sigma_{*}\end{cases}$ (3.25)
@@
$\tau_{\rm eq}$ $\sim\tau_{\rm eq}^{G}\alpha_{\rm SD}^{-2}\frac{\ln\Lambda_{G}}{\ln\Lambda_{Y}}\times\begin{cases}\frac{\left(1-\frac{\alpha_{\rm SD}^{\rm(eff)}\alpha_{\rm SD}}{\alpha_{\rm SS}^{\rm(max)}}\right)^{3/2}}{\sqrt{1-\alpha_{\rm SD}^{\rm(eff)}}}&\sigma_{\chi}\gtrsim\sigma_{*}\\ {\left(\frac{\sigma_{*,0}}{\sigma_{\chi,0}}\right)}^{3/2}\left(1-\alpha_{\rm SD}^{\rm(eff)}\right)&\sigma_{\chi}\lesssim\sigma_{*}\end{cases}$ (3.26)